25. An equilibrium mixture of PCl5(g), PCl3(g), and Cl2(g) has partial pressures of 217.0 Torr, 13.2 Torr, and 13.2 Torr, respectively. A quantity of Cl2(g) is injected into the mixture, and the total pressure jumps to 263.0 Torr (at the moment of mixing). The system then re-equilibrates. The appropriate chemical equation is

Calculate the new partial pressures after equilibrium is reestablished. An equilibrium mixture of PCls(g), PCls(g), and Cl2(g) has partial pressures of 217.0 Torr, 13.2 Torr, and 13.2 Torr, respectively. A quantity of Cl2(g) is injected into the mixture, and the total pressure jumps to 263.0 Torr (at the moment of mixing). The system then re-equilibrates. The appropriate chemical equation is Calculate the new partial pressures after equilibrium is reestablished. Number PC,223.4 Torr Number 6.80 Torr Cl, Number PCT,26.4 Torr
Answer #1

Solution-

Here Kc = [217.0] / ( [13.2] [13.2] )

= 1.2454

[263.0 Torr] - [217.0 Torr ] - [13.2 Torr ]

= 32.8 torr Cl2 ( before equilibrium and after the addition)


[217.0 + x ] / ( [13.2 - x ] [32.8 - x] ) = 1.2454
Solving for x:
x = 6.40 torr

[PCl3] = 13.2 - 6.4 = 6.8 torr
[Cl2] = 32.8 - 6.4 = 26.4 torr
[PCl5] = 217.0 + 6.4 = 223.4 torr